# SAKURA: On the Multi-hop Reasoning of Large Audio-Language Models Based on Speech and Audio Information

### v1

@@ Abstract
To address this gap, we introduce SAKURA (Speech and Audio-based Question-answering Benchmark for Multi-hop Reasoning of Large Audio-Language Models), a novel benchmark for systematic evaluation of LALMs’ multi-hop reasoning using speech and audio information. SAKURA has four tracks based on speech/audio attributes including speaker gender, spoken language, speaker emotion, and animal sound. Each track includes single-hop and multi-hop sub-tracks, both comprising 500 multiple-choice questions. Single-hop questions require direct perception of an attribute (e.g., “What is the animal in the sound?”), whereas multi-hop questions demand further reasoning based on the extracted information (e.g., “Which of the listed features aligns best with the physical appearance of the animal in the sound?”). This design compels models to recall and integrate the extracted auditory cues when tackling multi-hop questions. Fig. 1 provides examples of single-hop and multi-hop reasoning paths. In total, SAKURA offers 4000 human-verified multiple-choice questions, providing not only a structured framework for evaluating LALMs’ speech/audio-based multi-hop reasoning but also a valuable resource for future research on their reasoning mechanisms.
@@
Overall, our contributions are: (1) introducing SAKURA, the first benchmark for systematically evaluating LALMs’ multi-hop reasoning, and (2) uncovering LALMs’ difficulties in integrating speech/audio for multi-hop reasoning. Relevant resources such as dataset, evaluation code, etc., can be found at https://github.com/b08202033/SAKURA.
@@ 2.1 Overview
In single-hop sub-tracks, models must extract an attribute directly from a speech/audio input (e.g., identifying the animal sound). In multi-hop sub-tracks, models must use the extracted information for further reasoning (e.g., determining a physical trait that matches the identified animal). Each sub-track contains 500 multiple-choice questions, each consisting of (1) a speech/audio input, (2) a textual question with a fixed number of choices, and (3) a golden answer. The multiple-choice format was adopted to facilitate standardized assessment and comparison. Additionally, speech/audio inputs are shared across sub-tracks within each track, enabling controlled evaluation of reasoning relative to perception accuracy. Fig. 1 provides a data example, and Table 1 shows key dataset statistics, showcasing its diversity without strong biases toward any particular ranges.
@@ 2.3 Textual question-answer pair generation
Fig. 1 illustrates our question-answer data generation pipeline. We used GPT-4o444gpt-4o-2024-11-20, https://openai.com to generate question-answer pairs. In the single-hop sub-tracks, questions directly queried target attributes. To enhance diversity, we used GPT-4o to paraphrase manually crafted questions, such as “What is the animal in the sound?”, producing multiple variations with differing phrasings and lengths. These paraphrased questions were then paired with speech/audio data, with the associated attribute label as the golden answer and distractors selected from alternative labels.
@@ 4.2 Fail to reason, or fail to reason with speech and audio?
To further investigate why LALMs struggle with multi-hop reasoning, we compared their performances under two conditions: (1) speech/audio-based reasoning, where models process raw speech/audio inputs, and (2) text-based reasoning, where these inputs are replaced by captions summarizing the relevant attribute information (e.g., “The animal making the sound is cat.”) alongside the original instructions (see Fig. 1). This comparison examines how LALMs reason using encoded auditory representations compared to an approximate textual equivalent summarizing relevant information. For each model, we filtered the multi-hop sub-tracks to include only instances where it correctly answered the corresponding single-hop questions. This ensured that failures in multi-hop reasoning stemmed from reasoning limitations rather than misperception of attributes.

### v2

@@ Abstract
To address this gap, we introduce SAKURA (Speech and Audio-based Question-answering Benchmark for Multi-hop Reasoning of Large Audio-Language Models), a novel benchmark for systematic evaluation of LALMs’ multi-hop reasoning using speech and audio information. SAKURA has four tracks based on speech/audio attributes including speaker gender, spoken language, speaker emotion, and animal sound. Each track includes single-hop and multi-hop sub-tracks, both comprising 500 multiple-choice questions. Single-hop questions require direct perception of an attribute (e.g., “What is the animal in the sound?”), whereas multi-hop questions demand further reasoning based on the extracted information (e.g., “Which of the listed features aligns best with the physical appearance of the animal in the sound?”). This design compels models to recall and integrate the extracted auditory cues when tackling multi-hop questions. Figure 1 provides examples of single-hop and multi-hop reasoning paths. In total, SAKURA offers 4000 human-verified multiple-choice questions, providing not only a structured framework for evaluating LALMs’ speech/audio-based multi-hop reasoning but also a valuable resource for future research on their reasoning mechanisms.
@@
Overall, our contributions are: (1) introducing SAKURA, the first benchmark for systematically evaluating LALMs’ multi-hop reasoning, and (2) uncovering LALMs’ difficulties in integrating speech/audio for multi-hop reasoning. Relevant resources such as dataset, evaluation code, etc., can be found at https://github.com/ckyang1124/SAKURA.
@@ 2.1 Overview
In single-hop sub-tracks, models must extract an attribute directly from a speech/audio input (e.g., identifying the animal sound). In multi-hop sub-tracks, models must use the extracted information for further reasoning (e.g., determining a physical trait that matches the identified animal). Each sub-track contains 500 multiple-choice questions, each consisting of (1) a speech/audio input, (2) a textual question with a fixed number of choices, and (3) a golden answer. The multiple-choice format was adopted to facilitate standardized assessment and comparison. Additionally, speech/audio inputs are shared across sub-tracks within each track, enabling controlled evaluation of reasoning relative to perception accuracy. Figure 1 provides a data example, and Table 1 shows key dataset statistics, showcasing its diversity without strong biases toward any particular ranges.
@@ 2.3 Textual question-answer pair generation
Figure 1 illustrates our question-answer data generation pipeline. We used GPT-4o444gpt-4o-2024-11-20, https://openai.com to generate question-answer pairs. In the single-hop sub-tracks, questions directly queried target attributes. To enhance diversity, we used GPT-4o to paraphrase manually crafted questions, such as “What is the animal in the sound?”, producing multiple variations with differing phrasings and lengths. These paraphrased questions were then paired with speech/audio data, with the associated attribute label as the golden answer and distractors selected from alternative labels.
@@ 4.2 Fail to reason, or fail to reason with speech and audio?
To further investigate why LALMs struggle with multi-hop reasoning, we compared their performances under two conditions: (1) speech/audio-based reasoning, where models process raw speech/audio inputs, and (2) text-based reasoning, where these inputs are replaced by captions summarizing the relevant attribute information (e.g., “The animal making the sound is cat.”) alongside the original instructions (see Figure 1). This comparison examines how LALMs reason using encoded auditory representations compared to an approximate textual equivalent summarizing relevant information. For each model, we filtered the multi-hop sub-tracks to include only instances where it correctly answered the corresponding single-hop questions. This ensured that failures in multi-hop reasoning stemmed from reasoning limitations rather than misperception of attributes.